# Zak-OTFS with Interleaved Pilots to Extend the Region of Predictable Operation

### v1

@@ I Introduction
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler(DD) domain, that is a quasi-periodic localized function defined by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$ where $\tau_{p}\,\nu_{p}=1$. The time-domain (TD) realization of the carrier is a pulsone, that is a train of pulses modulated by a tone where adjacent pulses are spaced $\tau_{p}$ seconds apart. The frequency-domain (FD) realization of the carrier is a train of pulses in the frequency domain (FD) modulated by a FD sinusoid where adjacent pulses are spaced $\nu_{p}$ Hz apart. We have shown [1] that the Zak-OTFS input-output (I/O) relation is predictable [^1] Predictability implies that the channel response to an input DD pulse at any arbitrary discrete DD location $(k_{p},l_{p})$ can be predicted from the knowledge of the channel response to a DD pulse at some other location. and non-fading [^2] Consider the DD domain energy distribution of the channel response to an input DD pulse. The I/O relation is said to be non-fading if the energy distribution around the pulse is invariant of its location. when the delay period of the pulsone is greater than the delay spread of the channel and the Doppler period of the pulsone is greater than the Doppler spread of the channel. We refer to this condition as the crystallization condition. When the crystallization condition holds, the taps of the effective DD domain channel filter can simply be read off from the DD domain response to a single pilot carrier waveform, and the I/O relation can be reconstructed for a sampled system that operates under finite duration and bandwidth constraints. Section II describes the Zak-OTFS system model.
@@
Why Zak-OTFS rather than OFDM? Perhaps the most important reason is that 6G propagation environments are changing the balance between time-frequency methods focused on OFDM signal processing and delay-Doppler (DD) methods focused on Zak-OTFS signal processing. OFDM signals live on a coarse information grid (i.e., integer multiples of the sub-carrier spacing), and cyclic prefix/carrier spacing are designed to prevent inter carrier interference (ICI). When there is no ICI, equalization in OFDM is relatively simple once the I/O relation is acquired. However, acquisition of the Input/Output (I/O) relation in OFDM is non-trivial and model-dependent, and the interaction of the OFDM carrier with the channel varies in both TD and FD. By contrast Zak-OTFS signals live on a fine information grid. [^3] In Zak-OTFS, information symbols are carried by DD pulses having locations separated by integer multiples of the inverse bandwidth along the delay domain and separated by integer multiples of the inverse subframe duration along the Doppler domain. Since information carrying DD pulses are located on a finer grid, they interfere with each other resulting in inter-carrier interference due to which equalization is more involved. However, when the crystallization condition holds the I/O relation can be read off from the response to a single pilot signal. In this paper we focus on acquiring the I/O relation.
We consider the challenge of supporting two types of user with very different delay-Doppler characteristics without changing the delay and Doppler periods of the Zak-OTFS modulation. For simplicity, we focus on supporting a second user with delay spread $\tau_{max}$ at most $\tau_{p}/2$ and Doppler spread $2\nu_{max}$ satisfying $\nu_{p}<2\nu_{max}<2\nu_{p}$. We would be able to support this user by halving the delay period and doubling the Doppler period, but then we cannot choose a single delay and Doppler period to support both users.
@@ II System model
$y_{\mbox{\scriptsize{dd}}}^{w_{rx}}(\tau,\nu)$ $=$ $w_{rx}(\tau,\nu)*_{\sigma}y_{\mbox{\scriptsize{dd}}}(\tau,\nu).$ (11)
This filtered signal is then sampled on the information lattice $\Lambda_{\mbox{\scriptsize{dd}}}$ resulting in the quasi-periodic discrete DD domain signal $y_{\mbox{\scriptsize{dd}}}[k,l]$. This discrete DD output signal is related to the input discrete DD signal $x_{\mbox{\scriptsize{dd}}}[k,l]$ through the input-output (I/O) relation [18, 1]
@@
where [^6] For any two discrete DD functions $a[k,l]$ and $b[k,l]$, the discrete twisted convolution between them i.e., $a[k,l]*_{\sigma}b[k,l]=\sum\limits_{k^{\prime},l^{\prime}\in{\mathbb{Z}}}a[k^{\prime},l^{\prime}]\,b[k-k^{\prime},l-l^{\prime}]\,e^{j2\pi l^{\prime}\frac{(k-k^{\prime})}{MN}}$. $h_{\mbox{\scriptsize{eff}}}[k,l]$ is the effective DD domain channel filter and $n_{\mbox{\scriptsize{dd}}}[k,l]$ are the DD domain noise samples. Note that if $h_{\mbox{\scriptsize{eff}}}k,l]$ is simply
@@
The guard region ${\mathcal{G}}$ separates the pilot region from the data region comprising locations $(k_{0},l_{0})\in{\mathcal{I}}$. Fig. 2 shows pilot, guard and data regions as strips in the Zak-OTFS subframe that run parallel to the Doppler axis.
@@ III Two Interleaved Pilots
For simplicity, we suppose $2\nu_{p}>2\nu_{max}\geq\nu_{p}$. We transmit two interleaved pilots at locations $(k_{p_{1}},0)$ and $(k_{p_{2}},0)$ in ${\mathcal{D}}$, and Fig. 4 illustrates how we surround each pilot with pilot and guard regions. Data is not transmitted in the pilot and guard regions. Fig. 5 illustrates the channel responses to the impulses forming the two pilots. For $i=1,2$, the response $h_{0,0}^{(i)}[k,l]$ to the impulse at $(k_{p_{i}},0)$ is shown in blue, and the response $h_{0,1}^{(i)}[k,l]$ to the impulse at $(k_{p_{i}},N)$ is shown in red. Since $2\nu_{p}>2\nu_{max}\geq\nu_{p}$, it is only the response of adjacent impulses that overlap along the Doppler axis.
@@
Interleaved pilots make it possible to change the aspect ratio of the crystallization condition without changing the fundamental periods, $\tau_{p}$ and $\nu_{p}$. Here, we have illustrated the method of interleaved pilots for the case $2\nu_{p}>\nu_{max}\geq\nu_{p}$. By symmetry, a similar method applies to the case $2\tau_{p}>\tau_{max}\geq\tau_{p}$, with the only difference being that the pilots are interleaved along the Doppler axis and are also multiplied with distinct known unit modulus complex scalars. This is required, since the pilot is periodic along the Doppler axis and therefore without any multiplication with complex scalars the corresponding equations (26) and (27) will not be linearly independent.
@@ IV Multiple Interleaved Pilots
The generalization to $Q$ interleaved pilots is required when $(Q-1)\nu_{p}\leq 2\nu_{max}<Q\nu_{p}$. We transmit $Q$ interleaved pilots at locations $(k_{p_{i}},0)\in{\mathcal{D}},i=1,2\cdots,Q$, separated by pilots regions ${\mathcal{P}}_{i}$ and guard regions. The response $y_{\mbox{\scriptsize{dd}}}^{(i)}[k,l]$ to the pilot $(k_{p_{i}},0)$ received in ${\mathcal{P}}_{i}$ is a linear combination of $Q$ distinct taps of $h_{\mbox{\scriptsize{eff}}}[k,l]$. The responses $y_{\mbox{\scriptsize{dd}}}^{(i)}[k,l]$, $i=1,2,\cdots,Q$, yield $Q$ linear equations in the $Q$ unknown taps, and since the pilot locations are distinct, these equations are linearly independent. We estimate the taps $h_{\mbox{\scriptsize{eff}}}[k,l]$, where $k=-1,\cdots,k_{max}$ with $l=-\frac{Q}{2}N,\cdots,0,\cdots,(\frac{Q}{2}-1)N$ if $Q$ is even, and $l=-\frac{(Q-1)}{2}N,\cdots,0,\cdots,\frac{(Q-1)}{2}N$ if $Q$ is odd.
@@ V Auto-ambiguity of Interleaved Pilots
Figs. 6, 7, 8 and 9 illustrate the effectiveness of two interleaved pilots through heatmaps of the auto-ambiguity and cross-ambiguity functions.
@@
Finally we analyze how estimation accuracy of the proposed linear estimation in Section III and Section IV depends on the spacing between the interleaved pilots. For simplicity, we take $Q=2$ and apply linear estimation. The determinant $\Delta$ of the $2\times 2$ linear system satisfies $|\Delta|=\frac{E_{p}}{2}\left|e^{-j2\pi\frac{k_{p_{1}}}{M}}\,-\,e^{-j2\pi\frac{k_{p_{2}}}{M}}\right|$. As the minimum pilot spacing $\min\left(\left|k_{p_{1}}-k_{p_{2}}\right|\,,\,M-\left|k_{p_{1}}-k_{p_{2}}\right|\right)$ decreases, the determinant $\Delta$ approaches 00, and the $2\times 2$ system becomes highly ill-conditioned. We expect estimation accuracy to degrade as the minimum pilot spacing decreases.

### v2

@@ I Introduction
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler(DD) domain, that is a quasi-periodic localized function defined by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$ where $\tau_{p}\,\nu_{p}=1$. The time-domain (TD) realization of the carrier is a pulsone, that is a train of pulses modulated by a tone where adjacent pulses are spaced $\tau_{p}$ seconds apart. The frequency-domain (FD) realization of the carrier is a train of pulses in the frequency domain (FD) modulated by a FD sinusoid where adjacent pulses are spaced $\nu_{p}$ Hz apart. We have shown [1] that the Zak-OTFS input-output (I/O) relation is predictable [^1] Predictability implies that the channel response to an input DD pulse at any arbitrary discrete DD location $(k_{p},l_{p})$ can be predicted from the knowledge of the channel response to a DD pulse at some other location. and non-fading [^2] Consider the DD domain energy distribution of the channel response to an input DD pulse. The I/O relation is said to be non-fading if the energy distribution around the pulse is invariant of its location. when the delay period of the pulsone is greater than the delay spread of the channel and the Doppler period of the pulsone is greater than the Doppler spread of the channel [2]. We refer to this condition as the crystallization condition. When the crystallization condition holds, the taps of the effective DD domain channel filter can simply be read off from the DD domain response to a single pilot carrier waveform, and the I/O relation can be reconstructed for a sampled system that operates under finite duration and bandwidth constraints. Section II describes the Zak-OTFS system model.
@@
Why Zak-OTFS rather than OFDM? Perhaps the most important reason is that 6G propagation environments are changing the balance between time-frequency methods focused on OFDM signal processing and delay-Doppler (DD) methods focused on Zak-OTFS signal processing. OFDM signals live on a coarse information grid (i.e., integer multiples of the sub-carrier spacing), and cyclic prefix/carrier spacing are designed to prevent inter carrier interference (ICI). When there is no ICI, equalization in OFDM is relatively simple once the I/O relation is acquired. However, acquisition of the Input/Output (I/O) relation in OFDM is non-trivial and model-dependent, and the interaction of the OFDM carrier with the channel varies in both TD and FD. By contrast Zak-OTFS signals live on a fine information grid. [^3] In Zak-OTFS, information symbols are carried by DD pulses having locations separated by integer multiples of the inverse bandwidth along the delay domain and separated by integer multiples of the inverse subframe duration along the Doppler domain. Since information carrying DD pulses are located on a finer grid, they interfere with each other resulting in inter-carrier interference due to which equalization is more involved. However, when the crystallization condition holds the I/O relation can be read off from the response to a single pilot signal. In this paper we consider acquiring the I/O relation.
Figure 1: Orthogonal/non-overlapping allocation of time-frequency (TF) resource to different users having distinct delay-Doppler profile.
In a time and bandwidth limited communication system, users can be allocated different non-overlapping time-frequency (TF) resources by simply shifting each user’s signal in time and frequency (orthogonal multiple-access). In this paper, we consider Zak-OTFS based orthogonal TF access (there is no multi-user interference). A typical orthogonal allocation for four user terminals (UTs) is illustrated in Fig. 1. One important issue with OFDM, is that its numerology (sub-carrier spacing) is not flexible. In OFDM, a single user subject to inter-carrier interference causes carrier spacing to increase for all users, since OFDM does not allow for users to have different sub-carrier spacing. This increases the overall cyclic-prefix (CP) overhead substantially since a large sub-carrier spacing implies a small OFDM symbol duration whereas the CP duration is determined by the delay-spread of the channel.
On the other hand, Zak-OTFS numerology is more flexible. Each user can choose a different delay/Doppler period parameter $(\tau_{p},\nu_{p})$ depending on its own channel, i.e., a user which experiences a higher Doppler spread would choose a higher $\nu_{p}$ so that the crystallization condition holds (in Fig. 1, UT3 and UT4 experience higher Doppler spreads than UT1 and UT2). However, the implementation of such a system with different delay/Doppler period parameters for different users is challenging. Therefore, in this paper we consider supporting users with different delay-Doppler characteristics without changing the delay and Doppler periods of their Zak-OTFS modulation. We propose interleaved DD domain pilots, so that with $Q$ interleaved pilots it is possible to accurately acquire the I/O relation as long as the Doppler spread $2\nu_{max}$ is less than $Q\nu_{p}$. Although the associated pilot/guard overhead is higher for larger number of interleaved pilots, each user separately configures its own number of interleaved pilots depending on only its own channel Doppler spread.
@@ II System model
$y_{\mbox{\scriptsize{dd}}}^{w_{rx}}(\tau,\nu)$ $=$ $w_{rx}(\tau,\nu)*_{\sigma}y_{\mbox{\scriptsize{dd}}}(\tau,\nu)$ (11)
This filtered signal is then sampled on the information lattice $\Lambda_{\mbox{\scriptsize{dd}}}$ resulting in the quasi-periodic discrete DD domain signal $y_{\mbox{\scriptsize{dd}}}[k,l]$. This discrete DD output signal is related to the input discrete DD signal $x_{\mbox{\scriptsize{dd}}}[k,l]$ through the input-output (I/O) relation [^6] The relation between discrete and continuous I/O relation for delay-only channels and doubly-spread channels is similar and has been described in detail in [19] (see Tables-I, II and III and the related discussion in [19]). In delay-only channels, the discrete-time I/O relation forms the basis for practical implementation and this is also the case for doubly-spread channels where the discrete-DD domain I/O relation (see (12)) forms the basis for practical implementation. [19, 1]
@@
where [^7] For any two discrete DD functions $a[k,l]$ and $b[k,l]$, the discrete twisted convolution between them i.e., $a[k,l]*_{\sigma}b[k,l]=\sum\limits_{k^{\prime},l^{\prime}\in{\mathbb{Z}}}a[k^{\prime},l^{\prime}]\,b[k-k^{\prime},l-l^{\prime}]\,e^{j2\pi l^{\prime}\frac{(k-k^{\prime})}{MN}}$. $h_{\mbox{\scriptsize{eff}}}[k,l]$ is the effective DD domain channel filter and $n_{\mbox{\scriptsize{dd}}}[k,l]$ are the DD domain noise samples. Note that $h_{\mbox{\scriptsize{eff}}}k,l]$ is simply
@@
The guard region ${\mathcal{G}}$ separates the pilot region from the data region comprising locations $(k_{0},l_{0})\in{\mathcal{I}}$. Fig. 3 shows pilot, guard and data regions as strips in the Zak-OTFS subframe that run parallel to the Doppler axis. We use the pulsones in the pilot region (yellow strip) to acquire the taps of $h_{\mbox{\scriptsize{eff}}}[k,l]$, and therefore they do not carry information symbols. We do not transmit data in the entirety of the yellow strip so that we are able to acquire the I/O relation for a wide range of Doppler spreads (support of $h_{\mbox{\scriptsize{eff}}}[k,l]$ is shown as a pink ellipse).
The guard region on the left of the pilot region is bigger compared to that on the right. This is because, the channel path delays are positive and therefore the delay spread of the effective DD domain channel filter $h_{\mbox{\scriptsize{eff}}}[k,l]$ is asymmetric, i.e., more for positive delay tap values and less for negative delay tap values. The guard region on the left is required to minimize interference to the pilot region from information symbols in the data region to the left and the guard region on the right is required to minimize interference from the data region to the right. Due to the asymmetric delay spread of $h_{\mbox{\scriptsize{eff}}}[k,l]$, the symbols in the data region to the left can spread farther to the right when compared to the symbols in the data region to the right which can spread only a few taps to the left.
For the single pilot in Fig. 3, the delay axis width of the guard region to the left of the pilot region is $k_{max}$ taps and that of the guard region to the right of the pilot region is one tap. The width of the pilot region is $(k_{max}+2)$ taps. Since no information symbols are transmitted in the guard and the pilot regions, the fractional pilot overhead (i.e., ratio of the number of pulsones which do not carry information to the total number of pulsones) is $(2k_{max}+3)/M$.
@@ III Two Interleaved Pilots
For simplicity, we suppose $2\nu_{p}>2\nu_{max}\geq\nu_{p}$. We transmit two interleaved pilots at locations $(k_{p_{1}},0)$ and $(k_{p_{2}},0)$ in ${\mathcal{D}}$, and Fig. 5 illustrates how we surround each pilot with pilot and guard regions. The purpose of data, guard and pilots regions is the same as for the single pilot Zak-OTFS frame in Fig. 3. Data is not transmitted in the pilot and guard regions. Just as in Fig. 3, in Fig. 5 also the pilot region (yellow strip) is used to acquire $h_{\mbox{\scriptsize{eff}}}[k,l]$.
Fig. 6 illustrates the channel responses to the impulses forming the two pilots. For $i=1,2$, the response $h_{0,0}^{(i)}[k,l]$ to the impulse at $(k_{p_{i}},0)$ is shown in blue, and the response $h_{0,1}^{(i)}[k,l]$ to the impulse at $(k_{p_{i}},N)$ is shown in red. Since $2\nu_{p}>2\nu_{max}\geq\nu_{p}$, it is only the response of adjacent impulses that overlap along the Doppler axis.
@@
for $-1\leq k\leq k_{max}$, $l=0,1,\cdots,N-1$. When $k_{p_{1}}\neq k_{p_{2}}$ (mod $M$), equations (26) and (27) are linearly independent, and it is possible to solve for $h_{\mbox{\scriptsize{eff}}}[k,l]$ and $h_{\mbox{\scriptsize{eff}}}[k,l-N]$. [^11] In general, the received samples $y_{\mbox{\scriptsize{dd}}}^{(1)}[k+k_{p_{1}},l]$ and $y_{\mbox{\scriptsize{dd}}}^{(2)}[k+k_{p_{2}},l]$ also contain noise and interference from data symbols. Therefore, for the simulations reported in Section VI, we consider the least squares (LS) estimate of the channel tap coefficients $h_{\mbox{\scriptsize{eff}}}[k,l]$ and $h_{\mbox{\scriptsize{eff}}}[k,l-N]$. This also holds for the general case of $Q$ interleaved pilots discussed in Section IV, where we need to estimate $Q$ channel tap coefficients from $Q$ linear equations with additive noise and data interference. The least squares (LS) estimate of $(h_{\mbox{\scriptsize{eff}}}[k,l],h_{\mbox{\scriptsize{eff}}}[k,l-N])^{T}$ is given by (LABEL:eqn72443) (see top of this page).
@@
For the Zak-OTFS frame with two interleaved pilots in Fig. 5, the fractional pilot overhead (i.e., ratio of the number of pulsones which do not carry information to the total number of pulsones) is double of that for the single pilot Zak-OTFS frame in Fig. 3, i.e., $2(2k_{max}+3)/M$.
@@
Interleaved pilots make it possible to change the aspect ratio of the crystallization condition without changing the fundamental periods, $\tau_{p}$ and $\nu_{p}$. Here, we have illustrated the method of interleaved pilots for the case $2\nu_{p}>\nu_{max}\geq\nu_{p}$. By symmetry, a similar method applies when we have channel scenarios with large delay spread, larger than $\tau_{p}$. For example, with $2\tau_{p}>\tau_{max}\geq\tau_{p}$, we apply the same method with the only difference being that the pilots are interleaved along the Doppler axis (instead of the delay axis) and are also multiplied with distinct known unit modulus complex scalars. This is required, since the pilot is periodic along the Doppler axis and therefore without any multiplication with complex scalars the corresponding equations (26) and (27) will not be linearly independent. Note that there are important wireless markets which have large cells.
@@ IV Multiple Interleaved Pilots
The generalization to $Q$ interleaved pilots is required when $(Q-1)\nu_{p}\leq 2\nu_{max}<Q\nu_{p}$. We transmit $Q$ interleaved pilots at locations $(k_{p_{i}},0)\in{\mathcal{D}},i=1,2\cdots,Q$, separated by pilots regions ${\mathcal{P}}_{i}$ and guard regions. The response $y_{\mbox{\scriptsize{dd}}}^{(i)}[k,l]$ to the pilot $(k_{p_{i}},0)$ received in ${\mathcal{P}}_{i}$ is a linear combination of $Q$ distinct taps of $h_{\mbox{\scriptsize{eff}}}[k,l]$. The responses $y_{\mbox{\scriptsize{dd}}}^{(i)}[k,l]$, $i=1,2,\cdots,Q$, yield $Q$ linear equations in the $Q$ unknown taps, and since the pilot locations are distinct, these equations are linearly independent. We estimate the taps $h_{\mbox{\scriptsize{eff}}}[k,l]$, where $k=-1,\cdots,k_{max}$ with $l=-\frac{Q}{2}N,\cdots,0,\cdots,(\frac{Q}{2}-1)N$ if $Q$ is even, and $l=-\frac{(Q-1)}{2}N,\cdots,0,\cdots,\frac{(Q-1)}{2}N$ if $Q$ is odd. With $Q$ interleaved pilots and guard regions on both sides (just as in Fig. 3 and Fig. 5 for single and two pilots), the fractional pilot overhead is $Q$ times that for the single pilot, i.e., $Q(2k_{max}+3)/M$.
@@ V Auto-ambiguity of Interleaved Pilots
In the Zak-OTFS frame for two interleaved pilots (see Fig. 5), one can choose the two pilot locations such that there is no data region in between enabling the use of a common guard region which can reduce overhead. However, the two pilot locations will then be closer than $M/2$ delay taps due to which the auto-ambiguity function of the interleaved pilot signal will have a period of only $N$ along the Doppler axis and not $2N$ due to which the I/O relation cannot be acquired for $\nu_{p}<2\nu_{max}<2\nu_{p}$ resulting in poor error rate performance.
We explain this in detail with the help of Figs. 7, 8, 9 and 10. We consider $M=64,N=24$. Fig. 7 shows the heatmap of the auto-ambiguity function of the transmitted two interleaved pilots $(Q=2)$ when the delay domain spacing $|k_{p_{2}}-k_{p_{1}}|=M/2=32$. It is observed that the auto-ambiguity function is supported on a rectangular lattice with periods $2N$ and $M/2$ respectively along the Doppler and delay axis as discussed above (based on (LABEL:eqn0937745)). However, when the delay domain spacing is only $7$ delay taps (as in Fig. 9, the auto-ambiguity function is no more supported on a lattice and repeats with a period of $N$ (and not $2N$) along the Doppler axis.
In Fig. 8 and Fig. 10, we plot the heatmap of the cross-ambiguity between the received and the transmitted two interleaved pilots, when their spacing is $32$ and $7$ respectively. We consider a six-path Veh-A channel [25] with $\nu_{max}=6$ KHz and $\nu_{p}=7.5$ KHz. Note that the Doppler spread is $122\nu_{max}=$ 12 KHz which is more than $\nu_{p}$ but less than $2\nu_{p}$ and this is why the taps of $h_{\mbox{\scriptsize{eff}}}[k,l]$ can be read-off accurately from the received samples within the red rectangle shown in Fig. 8, when the delay spacing between the pilots is $M/2=32$ taps. However, when the pilot spacing is $27\neq M/$ 2, the cross-ambiguity in Fig. 10 looks completely different due to aliasing of responses $h_{0,0}^{(i)}[k,l]$ and $h_{0,1}^{(i)}[k,l]$ due to which accurate acquisition of the effective channel $h_{\mbox{\scriptsize{eff}}}[k,l]$ is not possible.
@@
Next, we analyze how estimation accuracy of the proposed linear estimation in Section III and Section IV depends on the spacing between the interleaved pilots. For simplicity, we take $Q=2$ and apply linear estimation. The determinant $\Delta$ of the $2\times 2$ linear system satisfies $|\Delta|=\frac{E_{p}}{2}\left|e^{-j2\pi\frac{k_{p_{1}}}{M}}\,-\,e^{-j2\pi\frac{k_{p_{2}}}{M}}\right|$. As the minimum pilot spacing $\min\left(\left|k_{p_{1}}-k_{p_{2}}\right|\,,\,M-\left|k_{p_{1}}-k_{p_{2}}\right|\right)$ decreases, the determinant $\Delta$ approaches 00, and the $2\times 2$ system becomes highly ill-conditioned. We expect estimation accuracy to degrade as the minimum pilot spacing decreases.
The number of interleaved pilots that a user needs for reliable communication depends on the Doppler spread of its channel. In general, if the Doppler spread $2\nu_{max}$ lies between $(Q-1)\nu_{p}$ and $Q\nu_{p}$, then $Q$ interleaved pilots are required (see discussion in Section IV). With $Q$ interleaved pilots regularly spaced apart by $M/Q$ delay axis taps, the auto-ambiguity function of the transmitted $Q$ interleaved pilots is supported on a DD domain lattice having delay and Doppler period $M/Q$ and $NQ$ respectively. Since the lattice period along the Doppler axis is now $Q$ times greater than the regular lattice period $N$ (with a single pilot), Doppler domain aliasing does not happen as long as the Doppler spread is less than $Q\nu_{p}$. Therefore, each user will use a different number of interleaved pilots depending on the maximum Doppler spread it experiences. The above result is shown in the following theorem.
Theorem 1
With $Q$ interleaved pilots and $k_{p_{i}}=(i-1)M/Q$, $i=1,2,\cdots,Q$, the auto-ambiguity function $A_{x_{p},x_{p}}[k,l]$ in (33) is supported on the DD domain lattice $\Lambda_{Q}$ given by
$\Lambda_{Q}$ $=$ ${\Big{\{}}(k,l)\,{\Big{|}}\,k=n\frac{M}{Q},l=mQN,\,(n,m)\in{\mathbb{Z}}{\Big{\}}}.$ (38)
Proof:
See Appendix A. ∎